OpenStudy (anonymous):

List all the real and imaginary zeros of f(x) = x^3+x^2+ 13x - 15... 1 is a zero of the function

OpenStudy (anonymous):

did you factor? you already knew 1 was a zero, from the last question

OpenStudy (callisto):

I think the question asks to solve f(x) =0 Since the question gives you 1 is a zero of the function. (x-1) is a factor. Now, do the division to get the other factor first, can you?

OpenStudy (anonymous):

once you factor, you find the other zeros using the quadratic formula

OpenStudy (anonymous):

You said (x−1)(x^2+2x+15) This gives zeros of 1, 3, 5, 15?

OpenStudy (callisto):

Nope!~ f(x) =0 x^3+x^2+ 13x - 15 =0 (x−1)(x^2+2x+15) =0 So, x-1 =0 -(1) or (x^2+2x+15) =0 -(2) You can solve (1) easily. And solve (2) using quadratic formula!~

OpenStudy (callisto):

For a quadratic equation ax^2 + bx + c=0 \[x=\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]^ Quadratic formula

OpenStudy (anonymous):

i get \[(-4\sqrt{15}) /2\]

OpenStudy (callisto):

It doesn't seems right... \[x=\frac{-2 \pm \sqrt{2^2 - 4(1)(15)}}{2(1)} =?\]

OpenStudy (anonymous):

-2i√14?

OpenStudy (callisto):

Nope.. \[x=\frac{-2\pm \sqrt {-56}}{2} = \frac{-2\pm -2\sqrt {-14}}{2} = -1 \pm -14i\]

OpenStudy (anonymous):

\[\sqrt{14}i\]

OpenStudy (anonymous):

a typo

OpenStudy (callisto):

Sorry :S \[-1\pm -\sqrt{14}i\]

OpenStudy (anonymous):

Im still not sure how I get the real and imaginary zeros off that...

OpenStudy (callisto):

the real one is 1 the imaginary ones are \[-1\pm -\sqrt{14}i\] Actually, the question just asks to you to solve f(x) =0 ... I think..

OpenStudy (anonymous):

Thankssss!

OpenStudy (callisto):

Welcome!!~~~